OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (4, –2) and has a slope of –3?

OpenStudy (anonymous):

you have to use the formula (y-y1) = m(x-x1) where (x1,y1) are given points

OpenStudy (anonymous):

y+3x=10

OpenStudy (aroub):

y=mx+b , m=-3 , (4,-2) -2=-3(4)+b -2=-12+b 12-2=b 10=b so the equation is y=-3x+10

OpenStudy (anonymous):

Start with point slope form and your info. y-(-2)=-3(x-4)...................simplify y+2 = -3x+12...................put in slope intercept form. Solve for y. y=-3x+10.........................move x term to other side. 3x+y=10..........................be sure there are no fractions in the problem.